OpenStudy (anonymous):

simplify (4n(n+1)(2n+1)/6)+n

OpenStudy (anonymous):

you may use foil method on (n+1)(2n+1) first.

OpenStudy (anonymous):

Check it once more...

OpenStudy (anonymous):

(4(2n^2+3n+1)/6)+n

OpenStudy (anonymous):

4?? or 4n???

OpenStudy (anonymous):

It is 4n outside the brackets: You have only solved for (n+1)(2n+1)

OpenStudy (anonymous):

(4(2n^2+3n+1)/6)+n wrong.. it is 4n outside.. Yes or no??

OpenStudy (anonymous):

See go step by step tell me what do you get: (n+1)(2n+1) = ?? Solve this first..

OpenStudy (anonymous):

2n^2+3n+1

OpenStudy (anonymous):

Yes... Now do this: \(4n(2n^2 + 3n + 1)\) = ??

OpenStudy (anonymous):

You are right don't mind.. My mind is mad..

OpenStudy (anonymous):

Yes you are right dear... Now solve that what I have written there..

OpenStudy (anonymous):

8n^3+12n^2+4n

OpenStudy (anonymous):

Yes so 4n not 4.. now write the entire what you get..

OpenStudy (anonymous):

can I write that as (8n^3/6)+(12n^2/6)+(4n/6)

OpenStudy (anonymous):

\[\frac{8n^3 + 12n^2 + 4n}{6} + n \implies \frac{8n^3 + 12n^2 + 4n + 6n}{6}\] Solve this..

OpenStudy (anonymous):

Are you getting this step??? You can also write what you did above that is also correct.. It is all about taking the LCD..

OpenStudy (anonymous):

right

OpenStudy (anonymous):

(4n^3/)3+(2n^2)+(5n/3)

OpenStudy (anonymous):

Absolutely right dear...

OpenStudy (anonymous):

Or you can leave the result as such like this there is nothing wrong in this: \[\implies \frac{8n^3 + 12n^2 + 10n}{6} \implies \frac{2n^3 + 6n^2 + 5n}{3}\]

OpenStudy (anonymous):

thanks